# FairQuant: Certifying and Quantifying Fairness of Deep Neural Networks

### v1

@@ I Introduction
Our work focuses on individual fairness which, at a high level, requires that similar individuals are treated similarly [5]. Here, similar individuals are those who differ only in some legally protected input attribute (e.g., gender or race) but agree in the unprotected attributes [^1] This notion can be understood as counterfactual fairness [6] explained in more detail in Section IX. and being treated similarly means that the DNN generates the same classification output. Let the DNN be a function $f:X\rightarrow Y$ from input domain $X$ to output range $Y$, where an individual $x\in X$ is an input and a class label $y\in Y$ is the output. Assume that each input $x=\langle x_{1},\ldots,x_{D}\rangle$ is a $D$-dimensional vector, and $x_{j}$, where $1\leq j\leq D$, is a protected attribute. We say that the DNN is provably fair (certified) for the entire input domain $X$ if $f(x)=f(x^{\prime})$ holds for any two individuals $x\in X$ and $x^{\prime}\in X$ that differ only in $x_{j}$ but agree in the unprotected attributes ($\forall x_{i}$ where $i\neq j$). Conversely, the DNN is provably unfair (falsified) for input domain $X$ if $f(x)\neq f(x^{\prime})$ holds for any two individuals ($x\in X$ and $x^{\prime}\in X$) that differ only in the protected attribute. If the DNN is neither certified nor falsified, it remains undecided.
@@ 6       $result\leftarrow$ Fair
7else if  $(O_{lb}>0\wedge O^{\prime}_{ub}<0)\vee(O_{ub}<0\wedge O^{\prime}_{lb}<0)$  then
@@ VIII-A Benchmarks
We evaluate our method using three legally-protected input attributes. For Bank and German, we use age; for Adult, we use gender; and for Compas, we use race.444 For Bank and German, we use binarized age attribute provided by Fairify. For Compas, we binarize race attribute as {white, non-white} as in [25]. These are consistent with Fairify and other prior works in the fairness research.
@@ VIII-B Experimental Setup
We ran all experiments on a computer with 2 CPU, 4GB memory, and Ubuntu 16.04 Linux operating system. We set a time limit of 30 minutes for each DNN model. Our experiments were designed to answer three research questions:
@@
Fairify requires a parameter MS (maximum size of an input attribute) based on which it creates a fixed number of input partitions prior to certification. On the DNN models trained for Bank, German, and Adult, we used the default MS values (100, 100, and 10) for Fairify to create 510, 201, and 16000 partitions, respectively.
On the new DNN models trained for Compas, we set MS to a small value of 2 to create 20 partitions for Fairify. This was done to maximize Fairify’s performance such that it does not “choke” in verifying each input partition. By default, Fairify uses 100 seconds as “soft timeout” for each input partition and uses 30 minutes as “hard timeout” for the entire DNN. This means that it takes at most 100 seconds to verify a single input partition, and if unsolved, it just moves to the next partition, until the entire 30 minutes runs out.
@@ VIII-C1 Results for RQ 1
While our method (FairQuant) found counterexamples for all 32 DNNs, Fairify found counterexamples for only 20 of the 32 models. In addition, it found counterexamples for only one of the 7 newly added models.
Moreover, Columns 5 and 11 show that, on models where both tools found counterexamples, the number of counterexamples found by FairQuant is often thousands of times more. For example, the largest number of counterexamples found by Fairify is 194 (for GC-3) but the large number of counterexamples found by FairQuant is 71,012 (for AC-5).
@@ VIII-D Summary
In general, a smaller refinement depth allows FairQuant to terminate quickly. During our experiments, FairQuant terminated after 0.29s and 1.24s for GC-4 and GC-5, respectively, compared to the 4 minutes and 30 minutes taken by Fairify and yet returned better results. In fact, for GC-5, Fairify spent 30 minutes but failed to find any counterexample.
If we increase the refinement depth, by increasing the two threshold values of FairQuant, its quantification results will get even better.
@@ IX Related Work
Our method is the first scalable method for certifying and quantifying individual fairness of a deep neural network, and it outperforms the most closely related prior work, Fairify [10], which is the only tool currently available for certifying individual fairness of a DNN. To the best of our knowledge, no other methods can match the accuracy, scalability, and functionality of our method.
Our method differs from existing techniques for verifying other types of individual fairness properties for neural networks. For example, techniques in [26] and [27] tackle a different type of fairness property proposed in [28] that allows some perturbation for the non-protected attributes. They are orthogonal to our certification and quantification techniques. Other techniques [29] locally verify a type of property known as $\varepsilon$-fairness which, similar to robustness, is defined using a given input $x$ and a small constant for perturbation.
Group fairness is yet another type of fairness property, which can be verified using probabilistic techniques [14, 15, 30]. The difference between individual fairness and group fairness is that, while individual fairness requires similar individuals to be treated similarly, group fairness requires similar demographic groups to be treated similarly. Furthermore, dependency fairness is studied in [31] and [32]. All of these fairness properties are different from one another in varying degrees.
There are other prior works related to fairness of machine learning models. For example, existing works like [28][33], and [6] proposed fairness verification techniques for other types of machine learning models, but they are not applicable to deep neural networks. Testing techniques can quickly detect fairness violations in machine learning models [34, 35, 11, 36, 12, 37], but it does not provide formal guarantee that is important for certain applications. There are also techniques for improving fairness of machine learning models [29, 38, 39, 40, 41, 42, 43, 44], but they differ significantly from our method, which focuses on certifying and quantifying fairness of existing DNN models.
At a high level, our method is related to the large number of robustness verifiers for deep neural networks based on interval analysis [7, 20, 8, 45], SMT solving [46, 47] and mixed-integer linear programming [48]. While these verifiers can decide if a model is robust against adversarial perturbation, as explained earlier in Section 2, they cannot directly certify individual fairness. Other neural network verifiers that deal with differential verification [49, 50, 51] or equivalence verification [52] are also different, since they evaluate over two networks instead of one network.
@@ X Conclusion
We have presented FairQuant, a scalable method for certifying and quantifying individual fairness of a deep neural network over the entire input domain. It relies on sound abstraction during symbolic forward analysis to improve scalability, and iterative refinement to improve accuracy. In addition to certifying fairness, it is able to quantify the degree of fairness by computing the percentages of inputs whose classification outputs can be certified as fair or falsified as unfair. We have evaluated the method on a large number of DNN models trained using four popular fairness research datasets. The experimental results show that the method significantly outperforms state-of-the-art techniques in terms of both accuracy and scalability, as well as the ability to quantify the degree of fairness.

### v2

@@ I Introduction
Our work focuses on individual fairness which, at a high level, requires that similar individuals are treated similarly [5]. Here, similar individuals are those who differ only in some legally protected input attribute (e.g., gender or race) but agree in the unprotected attributes, and being treated similarly means that the DNN generates the same classification output.111 This notion can be understood as causal fairness [6] or dependency fairness [7], which is a non-probabilistic form of counterfactual fairness [8]. Let the DNN be a function $f:X\rightarrow Y$ from input domain $X$ to output range $Y$, where an individual $x\in X$ is an input and a class label $y\in Y$ is the output. Assume that each input $x=\langle x_{1},\ldots,x_{D}\rangle$ is a $D$-dimensional vector, and $x_{j}$, where $1\leq j\leq D$, is a protected attribute. We say that the DNN is provably fair (certified) for the entire input domain $X$ if $f(x)=f(x^{\prime})$ holds for any two individuals $x\in X$ and $x^{\prime}\in X$ that differ only in $x_{j}$ but agree in the unprotected attributes ($\forall x_{i}$ where $i\neq j$). Conversely, the DNN is provably unfair (falsified) for input domain $X$ if $f(x)\neq f(x^{\prime})$ holds for any two individuals ($x\in X$ and $x^{\prime}\in X$) that differ only in the protected attribute. If the DNN is neither certified nor falsified, it remains undecided.
@@ 6       $result\leftarrow$ Fair
7else if  $(O_{lb}>0\wedge O^{\prime}_{ub}<0)\vee(O_{ub}<0\wedge O^{\prime}_{lb}>0)$  then
@@ VIII-A Benchmarks
We evaluate our method using three legally-protected input attributes. For Bank and German, we use age; for Adult, we use gender; and for Compas, we use race.444 For Bank and German, we use binarized age attribute provided by Fairify. For Compas, we binarize race attribute into {white, non-white} as done in [27]. These are consistent with Fairify and other prior works in the fairness research.
@@ VIII-B Experimental Setup
We ran all experiments on a computer with 2 CPU, 4GB memory, and Ubuntu 20.04 Linux operating system. We set a time limit of 30 minutes for each DNN model. Our experiments were designed to answer three research questions:
@@
Fairify requires a parameter MS (maximum size of an input attribute) based on which it creates a fixed number of input partitions prior to certification. On the DNN models trained for Bank, German, and Adult, we used the default MS values (100, 100, and 10) for Fairify to create 510, 201, and 16000 partitions, respectively. On the new DNN models trained for Compas, we set MS to a small value of 2 to create 20 partitions for Fairify. This was done to maximize Fairify’s performance such that it does not “choke” in verifying each input partition.
By default, Fairify uses 100 seconds as “soft timeout” for each input partition and uses 30 minutes as “hard timeout” for the entire DNN. This means that it takes at most 100 seconds to verify a single input partition, and if unsolved, it just moves to the next partition, until the entire 30 minutes runs out.
@@ VIII-C1 Results for RQ 1
While our method (FairQuant) found counterexamples for all 32 DNNs, Fairify found counterexamples for only 20 of the 32 models. In addition, it found counterexamples for only one of the 7 newly added models. Moreover, Columns 5 and 11 show that, on models where both tools found counterexamples, the number of counterexamples found by FairQuant is often thousands of times more. For example, the largest number of counterexamples found by Fairify is 194 (for GC-3) but the large number of counterexamples found by FairQuant is 71,012 (for AC-5).
@@ VIII-D Summary
In general, a smaller refinement depth allows FairQuant to terminate quickly. During our experiments, FairQuant terminated after 0.29s and 1.24s for GC-4 and GC-5, respectively, compared to the 4 minutes and 30 minutes taken by Fairify and yet returned better results. In fact, for GC-5, Fairify spent 30 minutes but failed to find any counterexample. If we increase the refinement depth, by increasing the two threshold values of FairQuant, its quantification results will get even better.
@@ IX Related Work
Our method is the first scalable method for certifying and quantifying individual fairness of a deep neural network, and it outperforms the most closely related prior work, Fairify [12]. To the best of our knowledge, no other methods can match the accuracy, scalability, and functionality of our method.
Our method differs from existing techniques for verifying individual fairness properties for neural networks. Libra [7, 28] uses abstract domains to perform verification but is limited in scalability due to the expensive pre-analysis, where a network of 20 hidden nodes takes several hours even with leveraging multiple CPUs. DeepGemini [29], which outperforms Libra, is built on top of Marabou [30], a SMT-based neural network verification tool, and thus shares the same limitations of Fairify. Furthermore, it only evaluates on networks with up to around 250 hidden neurons. Other works have tackled neural network verification of different definitions of individual fairness. Benussi et al. [31] and Khedr et al. [32] proposed different methods to certify a definition of global individual fairness proposed in [33]. Ruoss et al. [34] verify a type of a local individual fairness property that is similar to local robustness, given an input $x$ and a small constant $\varepsilon$ for perturbation. This is different from a global perspective we have discussed so far in this paper.
Group fairness is yet another type of fairness property, which can be verified using probabilistic techniques [16, 17, 35]. The difference between individual fairness and group fairness is that, while individual fairness requires similar individuals to be treated similarly, group fairness requires similar demographic groups to be treated similarly.
There are other prior works related to fairness verification of other types of machine learning models [33, 36, 8], but they are not applicable to deep neural networks. Testing techniques can quickly detect fairness violations in machine learning models including neural networks [37, 6, 13, 38, 14, 39], but it does not provide formal guarantee that is important for certain applications. There are also techniques for improving fairness of machine learning models [40, 41, 42, 43, 44, 45, 46], which are orthogonal to our method that focuses on certifying and quantifying fairness of existing DNN models.
At a high level, our method is related to the large number of robustness verifiers for deep neural networks based on interval analysis [9, 22, 10, 47], SMT solving [23, 48, 30, 49], and mixed-integer linear programming [50, 51]. While these verifiers can decide if a model is robust against adversarial perturbation, they cannot directly certify individual fairness, as explained earlier in Section II. Other neural network verifiers that deal with differential [52, 53, 54] or equivalence verification [55] are also different, since they evaluate over two networks instead of one network.
@@ X Conclusion
We have presented FairQuant, a scalable method for certifying and quantifying individual fairness of a deep neural network over the entire input domain. It relies on sound abstraction during symbolic forward analysis to improve scalability, and iterative refinement based on backward analysis to improve accuracy. In addition to certifying fairness, it is able to quantify the degree of fairness by computing the percentages of inputs whose classification outputs can be certified as fair or falsified as unfair. We have evaluated the method on a large number of DNN models trained using four popular fairness research datasets. The experimental results show that the method significantly outperforms state-of-the-art techniques in terms of both accuracy and scalability, as well as the ability to quantify the degree of fairness.